OpenStudy (anonymous):

A candy store makes a 6–pound mixture of gummy bears, jelly beans, and gobstoppers. The cost of gummy bears is $1.00 per pound, jelly beans cost $3.00 per pound, and gobstoppers cost $1.00 per pound. The mixture calls for three times as many gummy bears as jelly beans. The total cost of the mixture is $8.00. How much of each ingredient did the store use?

OpenStudy (anonymous):

Could you be more specific, @ericaneal1 ?

OpenStudy (anonymous):

ok 3 times 1 is 3 there you go keep doing that then add if you dont get the divideing way

OpenStudy (anonymous):

That sounds like it would take a really long time. I would recommend algebra. There are three unknowns, so you need three equations. Do you see the three separate pieces of information, @jacobblocker ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The first thing you want to do (after reading the problem twice to understand the question being asked, and to gather the necessary information), is to define your variables. I would do something like G = pounds of gummy bears, J = pounds of jelly beans, S = pounds of gobstoppers.

OpenStudy (anonymous):

Using those variables, set up an equation for each sentence in the word problem.

OpenStudy (anonymous):

is it something like this g+j+s=8 g=3j 1g+1j+3s=6

OpenStudy (anonymous):

Hmm, pretty close. The first statement is about weight, and the weight was 6 pounds, not 8. In the third statement, it is the jelly beans that cost 3 $/Lb. not the gobstoppers, and the total cost is 8, not 6$

OpenStudy (anonymous):

You got the right idea though. :-)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

So, ultimately you want to get down to one equation with one unknown in it, so you can solve for its value. Right now you have three equations, so you'll first want to combine those in a way to make two equations with two unknowns, then do some substitution to combine those into a single equation. My first hint is to substitute the G = 3J expression into the other two equations, then use those to solve for J and S.

OpenStudy (anonymous):

ill try to

OpenStudy (anonymous):

Yeah, keep playing around with it, and let me know how far you get. I'll work on it too and see if I can find a solution.

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

What do you have so far?

OpenStudy (anonymous):

i dont know how to eliminate the equations

OpenStudy (anonymous):

Did you make this substitution? \[\large G=(3J).\] \[\large (G)+J+S=6 \rightarrow (3J)+J+S=6.\] \[\large (G)+3J+S=8 \rightarrow (3J)+3J+S=8.\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do that first; it'll make your life a lot easier.

OpenStudy (anonymous):

well you need to do that first

OpenStudy (anonymous):

You'll have two equations in J and S; you can solve the first one for S in terms of J and substitute that into the second one to get a value for J, then it's back-substitution to find the remaining values.

OpenStudy (anonymous):

A solution using Mathematica's Minimize function is attached.